OpenStudy (anonymous):

can anyone simplifiy this (csc(θ) (1-cos(θ)^2 ))/cos(θ)

OpenStudy (larseighner):

\(1 - \cos^2\theta) = \sin^2 \theta \)

OpenStudy (larseighner):

\(\csc\theta = {1 \over {\sin\theta}}\)

OpenStudy (larseighner):

so the numerator is \( {1\over{\sin\theta}}(\sin^2\theta) = \sin\theta\)

OpenStudy (larseighner):

\( {{\sin\theta} \over {\cos\theta}} = \tan\theta \)

OpenStudy (anonymous):

thanks for that, I am a bit older than the average student and appreciate your reply